# LASER: Lip Landmark Assisted Speaker Detection for Robustness

### v1

@@ 3.2.1 Lip Landmark Track Encoding
Integrating discrete 2D coordinate information with continuous visual signals during training poses a challenge. To address this, we design a simple-yet-effective encoding function $\mathcal{T}(M)$ that converts the 2D discrete coordinate information to continuous 2D feature maps matching the spatial dimensions of the visual features. Specifically, the encoding process $h=\mathcal{T}(M)$ for each landmark ${(x_{i},y_{i})}$ is defined as follows:
@@
The encoding process transforms the “lip landmark track” into features $h\in\mathbb{R}^{T\times K\times 2\times H\times W}$, where $K$ is the total number of lip landmarks, and 2 corresponds to the $x$ and $y$ coordinates. We use the lightweight facial landmark detector from Mediapipe [16] to extract $K=82$ lip landmarks.
@@ 3.2.2 Lip Landmark Track Feature Aggregation
We concatenate the lip landmark track feature $h^{\prime}$ with the face track visual features output by a 3D convolution layer and forward the concatenated feature to a ResNet for frame-level processing. The resulting feature is then passed through a video temporal network [14] to obtain temporal visual representations $f_{v*}\in\mathbb{R}^{T\times D}$. In this way, the visual encoder as a whole not only incorporates cues from frame-level lip landmarks but also leverages temporal modeling to explicitly capture the dynamics of lip movement across video frames.
Our visual encoder can be integrated to state-of-the-art ASD models [26, 25, 15, 5] and can be trained end-to-end by the ASD training objectives described in Equation 1. As we shall see in the experiments, LASER enhances state-of-the-art ASD models across various evaluation protocols.
@@ 3.3 Addressing Missing Lip Landmark Tracks
While the proposed LASER achieves strong performance across various evaluation scenarios, we observe that the facial landmark detector does not always return lip landmarks for every face track in videos because of low resolution, occlusion, or extreme angles in the facial frames. For example, on the AVA-ActiveSpeaker dataset [18], the detector fails to predict lip landmarks for approximately 15% of face tracks. At test time, the absence of lip landmarks results in suboptimal visual features, which can negatively impact the model’s performance.
@@
where $j$ denotes the class index of prediction. We apply stop gradient operation on $f_{v*}$ (predictions made with lips track encoding) when computing the consistency loss so that gradients only flow through $f_{v}$ (predictions without lips track encoding), encouraging the model to match the predictions it makes with lips track encoding. This approach ensures the model learns to make reliable predictions even when mouth landmarks are unavailable at test time.
@@
where $\mathcal{L}_{\text{asd}}$ is the default training objective of ASD models [26, 25, 15, 5] from Section 3.1. By enforcing consistent predictions between face tracks with and without lip track encoding, our model achieves robust performance even on challenging face tracks where facial landmark detection may fail. Notably, as demonstrated in our experiments, this approach allows us to eliminate the reliance on the facial landmark detector altogether, using only face tracks at test time without sacrificing performance.
@@ 4 Experiments
In this section, we evaluate the effectiveness of our proposed Lip landmark Assisted Speaker detection for Robustness (LASER) method. We assess its performance on multiple datasets and compare it against baseline models to demonstrate improvements in active speaker detection tasks in different evaluation scenarios. We also ablate our design choices.
@@ 4.1.1 Implementation Details
To demonstrate the efficacy of LASER, we implement it with the state-of-the-art ASD model LoCoNet [26]. We closely follow the training details for fair comparison. Specifically, we use batch size of 4 and sample 200 frames per training example. We use Adam [6] as our optimizer with learning rate of $5\times 10^{-5}$ that reduces by 0.5% per epoch. Each model is trained with 25 epochs on 4 RTX 2080 GPUs. For ASD training objectives, we follow the LoCoNet [26] setting $\lambda_{av}=1$, $\lambda_{a}=0.4$, and $\lambda_{v}=0.4$. Random resizing, cropping, horizontal flipping, and rotations are used as visual data augmentation operations and a randomly selected audio signal from the training set is added as background noise to the target audio [25].
4.1.2 Datasets
We experiment on common Active Speaker Benchmark datasets AVA-ActivateSpeaker [18], Talkies [2] and ASW [12]. AVA-ActivateSpeaker is the current biggest dataset for the Active Speaker Detection task. It consists of 262 videos spanning over 38 hours with over 38,500 tracks and over 3.65 million faces. Talkies [2] is one of the in-the-wild active speaker datasets with more than 799,000 faces detected from over 10,000 unique videos. On average, there are about 2.3 faces per frame which is greater than the average of 1.6 faces per frame of AVA-ActiveSpeaker. ASW [12] is an in-the-wild active speaker dataset along side Talkies. It contains 212 videos from the VoxConverse dataset [4], and consists of a total of 11,551 face tracks spanning over 30.9 hours.
@@ 4.1.3 Evaluation Protocols
(1) Evaluation with synchronized audios: This is the standard evaluation protocol, where both training and testing data have synchronized audio and face tracks. We evaluate on both in-domain scenarios (where training and testing data come from the same dataset) and out-of-domain scenarios (where testing data comes from different datasets). We use AVA-ActiveSpeaker as our in-domain datasets and Talkies and ASW as our out-of-domain evaluation datasets and use mAP as the evaluation metric following common practice [18].
(2) Evaluation with unsynchronized audios: This protocol simulates a non-speaking scenario by introducing a short delay at the beginning of the audio track in the test data, creating a minor desynchronization with the face tracks. This setup allows us to evaluate the model’s sensitivity to audiovisual synchronization. We use the same dataset partition as (1) and report the average accuracy across all frames.222mAP is not a proper metrics for this scenario because the ground-truth of all example frames are negative (non-speaking).
@@
Figure 4: Evaluation with unsynchronized audios. We use the same datasets as the evaluation of synchronized audios and report the per-frame accuracy on in-domain datasets (AVA-ActiveSpeaker) and out-of-domain datasets (Talkies and ASW). LASER consistently outperforms LoCoNet [26] under this evaluation protocol.
4.1.4 Hyperparameter Selection
LASER introduces two key hyperparameters: the number of dense feature maps for lip track encoding, $S$, and the weight of the consistency loss, $\lambda_{c}$. We set $S=4$ and $\lambda_{c}=1$ for all evaluations. They are determined through a holdout data evaluation using unsynchronized audio where the audio was randomized by reversing the original track. This scenario differs from evaluation protocol (2), which involves a short delay in audio tracks. Our hyperparameter selection process avoids the risk of “tuning on the test set,” ensuring the generalizability of our method.
4.2 Main Results
@@
TalkNet [25] 92.3 - -
EASEE [3] 94.1 86.7 -
MAAS [2] 88.8 79.7 -
LoCoNet [26] 95.2 88.4 88.4
LASER (ours) 95.3 89.0 88.9
LoCoNet w/ CL [10] 95.5 88.3 88.5
LASER + w/ CL (ours) 95.4 89.7 89.5
@@
Table 1: Evaluation with synchronized audios. The models are trained on AVA-ActiveSpeaker dataset only and evaluated on AVA, as in-domain evaluation and Talkies and ASW as out-of-domain evaluations. CL denotes contrastive learning.
4.2.1 Evaluation with Synchronized Audios
@@
We first evaluate our LASER with synchronized audios where the model is trained on the training data of AVA-ActiveSpeaker and evaluated on validation set of AVA-ActiveSpeaker, Talkies, and ASW. As shown in Table 1 (top), LASER still improves upon LoCoNet for in-domain evaluations when the performance is already saturated on AVA. When evaluating on Talkies and ASW whose data distribution is different from AVA, LASER achieves 89.0 and 88.9 mAP respectively.
@@
Recent work [10] also proposes to add an auxiliary contrastive loss to LoCoNet which brings audio-visual embedding pairs closer when they originate from the same video frame, while pushing them apart otherwise. As shown in Table 1 (bottom), while the contrastive loss enhances LoCoNet’s in-domain performance by 0.3 mAP, it offers limited gains in out-of-domain scenarios. Interestingly, when the same contrastive loss is applied to LASER, our method achieves an additional 1.4 and 1.0 mAP improvement on the Talkies and ASW benchmarks, respectively. These results highlight both the effectiveness and general applicability of LASER.
Layer Regular Audio Noise Audio
3D CNN 92.6 86.6
ResNet 95.1 86.7
Table 2: Ablation study on lip track encoding integration. Integrating lip track encoding before 3D-CNN layer (i.e, on RGB frames) results in suboptimal performance.
4.2.2 Evaluation with Unsynchronized Audios
We further demonstrate the robustness of LASER by evaluating with unsynchronized audios. We add short delays of 0.2s to 3s at the beginning of the audio track to create non-speaking scenarios and evaluate the model with original the face track and the shifted audio track. As shown in Figure 4, our LASER consistently outperforms LoCoNet by a large margin across different amount of delays in the audio. These results demonstrate the robustness of LASER in response to unsynchronized audios.
@@ 4.2.2 Evaluation with Unsynchronized Audios
Consistency Loss Lip Track Encoding Regular Audio Noise Audio
95.1 86.7
95.3 87.5
95.3 87.5
Table 3: Ablation study on the consistency loss. Consistency loss improves the performance on AVA-ActiveSpeaker dataset with approximately 15% videos with missing lips landmark track. Furthermore, with consistency loss, the model achieves similar performance even without lip track encoding, eliminating the dependency of lips track at test time.
@@ 4.3 Ablation Studies
We present ablation studies of LASER in this section. For all results in this section, we use the AVA-ActiveSpeaker dataset in both training and testing. As stated before, we select all our hyperparameters based on the performance with noise audio tracks. We include the results for regular audio tracks in this section for reference only.
4.3.1 Integrating Lip Track Encoding to Visual Features

### v2

@@ 3.2.1 Lip Landmark Track Encoding
Integrating discrete 2D coordinate information with continuous visual signals during training poses a challenge. To address this, we design a simple-yet-effective encoding function $\mathcal{T}(M)$ that converts the 2D discrete coordinate to continuous 2D feature maps matching the spatial dimensions of the visual features. Specifically, the encoding process $h=\mathcal{T}(M)$ for each landmark $(x_{i},y_{i})$} is defined as follows:
@@
The encoding process transforms the “lip landmark track” into features $h\in\mathbb{R}^{T\times K\times 2\times H\times W}$, where $K$ is the number of lip landmarks, and 2 corresponds to the $x$ and $y$ coordinates. We use the lightweight facial landmark detector from Mediapipe [lugaresi2019mediapipe] to extract $K=82$ lip landmarks.
@@ 3.2.2 Lip Landmark Track Feature Aggregation
We concatenate the lip landmark track feature $h^{\prime}$ with face track visual features from a 3D convolution layer, passing the combined feature through a ResNet for frame-level processing. The output is then fed into a video temporal network [lea2016temporal] to generate temporal visual representations $f_{v*}\in\mathbb{R}^{T\times D}$. In this way, the visual encoder not only incorporates cues from frame-level lip landmarks but also leverages temporal modeling to explicitly capture the dynamics of lip movement across video frames.
Our visual encoder can be integrated to state-of-the-art ASD models [wang2024loconet, tao2021someone, liao2023light, datta2022asd] and can be trained end-to-end by the ASD training objectives described in Equation 1. As we shall see in the experiments, LASER enhances state-of-the-art ASD models across various evaluation protocols.
@@ 3.3 Addressing Missing Lip Landmark Tracks
While LASER achieves strong performance across various evaluation scenarios, we observe that the facial landmark detector does not always return lip landmarks for every face track because of low resolution, occlusion, or extreme angles in the facial frames. For example, our facial landmark detector fails to detect any landmarks on approximately 15% of the face track on the AVA-ActiveSpeaker dataset [roth2020ava]. At test time, the absence of lip landmarks results in suboptimal visual features, which can negatively impact the model’s performance.
@@
where $j$ denotes the class index of prediction. We apply stop gradient operation on $f_{v*}$ (predictions made with lips track encoding) when computing the consistency loss so that gradients only flow through $f_{v}$ (predictions without lips track encoding), encouraging the model to match the predictions it makes with lips track encoding, implicitly guiding the model to focus on the lips-audio synchronization even when the lips landmarks are not provided.
@@
where $\mathcal{L}_{\text{asd}}$ is the default training objective of ASD models [wang2024loconet, tao2021someone, liao2023light, datta2022asd] from Section 3.1. By enforcing consistent predictions between face tracks with and without lip track encoding, our model achieves robust performance even on challenging face tracks where facial landmark detection may fail. Notably, as demonstrated in our experiments, this approach allows us to completely eliminate the reliance on the facial landmark detector, using only face tracks at test time without sacrificing performance.
4 LASER-bench
Recall that LASER leverages explicit lip information to improve audio-visual correspondence, which is particularly valuable in acoustically noisy settings where background sounds may interfere with speech perception. However, existing ASD benchmarks offer limited support for evaluating this aspect. AVA-ActiveSpeaker [roth2020ava], though large in scale, consists mainly of clean audio from scripted movie scenes, while ASW [kim2021look] includes noisy, in-the-wild videos but remains small in size. To bridge this gap, we introduce LASER-bench, a large-scale benchmark designed to evaluate ASD robustness under authentic background noise.
Statistics AVA-Val [roth2020ava] Talkies-Val [alcazar2021maas] ASW-Val [kim2021look] LASER-bench (Ours)
Total face tracks 8K 6.8K 3.5K 4.9K
Total face crops 760K 235K 171K 738K
Table 1: Dataset comparison. We compare relevant statistics of MADE with three other well-known datasets for active speaker detection (left columns) and LASER-bench (right columns).
4.1 Data Curation.
In creating LASER-bench, we aim to assemble a diverse dataset that captures complex acoustic scenarios, including off-screen speech, crowd noise, and environmental noise. To achieve this goal, we primarily use YouTube to source our clips, focusing on modern videos that typically include spoken dialogue and multiple visible individuals. After gathering an initial pool of vlogs, street interviews, podcasts/interviews, and multi-person TV shows through relevant search terms, we then filter out any content involving sensitive or potentially objectionable topics. Then, we follow the procedure described in AVA-ActiveSpeaker [roth2020ava] by keeping the face tracks that are at least 0.2 seconds and interpolating to create missing detection inside the tracks. However, we do not limit the length of each track like AVA-ActiveSpeaker [roth2020ava], and every annotation is gone through two rounds of verification by our annotators.
4.2 Dataset Statistics
Table 1 provides a concise comparison of four active‐speaker benchmarks, listing total face tracks and face crops for each: AVA‐Val [roth2020ava] (8K tracks, 760K crops), Talkies‐Val [alcazar2021maas] (6.8K tracks, 235K crops), ASW‐Val [kim2021look] (3.5K tracks, 171K crops) and our LASER-bench (4.9K tracks, 738K crops). Notably, while ASW‐Val incorporates real background noise to simulate challenging acoustic conditions, LASER‐bench preserves that same noise realism at a much larger scale - offering over 1.4× more face tracks and more than 4× the number of face crops - thereby furnishing a richer resource for evaluating speaker detection models under noisy, in‐the‐wild scenarios.
@@ 5 Experiments
In this section, we evaluate the effectiveness of our proposed approach (LASER). We assess its performance on multiple datasets and compare it against baseline models to demonstrate improvements across different evaluation scenarios. We also ablate our design choices.
@@ 5.1.1 Implementation Details
We implement LASER with the ASD models including LoCoNet [wang2024loconet], TalkNet [tao2021someone], and Light-ASD [liao2023light]. We closely follow the training details of each model for fair comparison (details in Appendix). For ASD training objectives, we follow the LoCoNet and TalkNet, setting $\lambda_{av}=1$, $\lambda_{a}=0.4$, and $\lambda_{v}=0.4$. Random resizing, cropping, horizontal flipping, and rotations are used as visual data augmentation operations and a randomly selected audio signal from the training set is added as background noise to the target audio [tao2021someone]. For LASER, We set $S=4$ and $\lambda_{c}=1$ for all evaluations and ablation studies of these hyper-parameters can be found in Appendix.
@@ 5.1.2 Evaluation Protocols
(1) Standard Evaluation: We evaluate on both in-domain scenarios (where training and testing data come from the same dataset) and out-of-domain scenarios (where testing data comes from different datasets). We use AVA-ActiveSpeaker [roth2020ava] as our in-domain datasets and Talkies [alcazar2021maas] and ASW [kim2021look] as our out-of-domain evaluation datasets and use mAP as the evaluation metric following common practice [roth2020ava].
(2) Evaluation on LASER-bench. In addition to the synchronization schema, we evaluate LASER on LASER-bench using the same metric mentioned above. In particular, we divide LASER-bench into two subsets: Low Noise and High Noise. To do so, for each audio track, we first remove speech segments using a Voice Activity Detection (VAD) tool [SileroVAD], then compute the Root Mean Square (RMS) energy over the remaining background audio. A threshold of 0.03 RMS distinguishes low and high noise levels.
(3) Evaluation with Unsynchronized audios: This protocol simulates non-speaking scenarios where visual frames and audio tracks are misaligned, a critical test for applications like web conferencing. For instance, if a muted participant talks to someone else while the main speaker is active, the model should correctly classify the muted participant as “not speaking” to avoid confusion. We evaluate using two protocols: 1) replacing the original audio with another video’s audio and 2) introducing a short audio delay. The evaluation uses the same dataset in (1), with results reported as average accuracy across all frames. 222mAP is unsuitable here as all ground-truth frames are negative (non-speaking).
@@
Figure 3: Evaluation with unsynchronized audios. We use the same datasets as the evaluation of synchronized audios and report the per-frame accuracy on in-domain datasets (AVA-ActiveSpeaker) and out-of-domain datasets (Talkies and ASW). LASER consistently outperforms LoCoNet [wang2024loconet] under this evaluation protocol.
@@
EASEE [alcazar2022end] 94.1 86.7 -
MAAS [alcazar2021maas] 88.8 79.7 -
TalkNet [tao2021someone] 92.2 85.9 85.8
TalkNet + LASER 92.5 87.6 86.6
Light-ASD [liao2023light] 93.5 87.6 87.4
Light-ASD + LASER 93.8 88.1 87.6
LoCoNet [wang2024loconet] 95.2 88.4 88.4
LoCoNet + LASER 95.3 89.0 88.9
LoCoNet w/ CL [jung2024talknce] 95.5 88.3 88.5
LoCoNet w/ CL + LASER 95.4 89.7 89.5
@@
Table 2: Results of models trained on AVA-ActiveSpeaker [roth2020ava]. The models are trained on AVA-ActiveSpeaker dataset only and evaluated on AVA, as in-domain evaluation and Talkies and ASW as out-of-domain evaluations (underlined). CL denotes contrastive learning. Results are reproduced by the methods’ codebase.
5.2 Standard Evaluation
We further evaluate LASER on AVA-ActiveSpeaker, the most widely used benchmark for active speaker detection. Training is conducted on the AVA training set, with in-domain evaluations on the AVA-val set and out-of-domain evaluations on Talkies [alcazar2021maas] and ASW [kim2021look], following LoCoNet [wang2024loconet].
@@ 5.2 Standard Evaluation
As shown in Table 2, LASER still improves upon LoCoNet for in-domain evaluations when the performance is already saturated on AVA. When evaluating on Talkies and ASW whose data distribution is different from AVA, LASER achieves 89.0 and 88.9 mAP respectively. Moreover, for efficient architectures like TalkNet or Light-ASD, LASER advances the performance on all three benchmarks. For example, LASER improves TalkNet’s evaluation on Talkies and ASW by 1.7 and 0.8 mAP respectively.
@@
Recent work [jung2024talknce] also proposes to add an auxiliary contrastive loss to LoCoNet which brings audio-visual embedding pairs closer when they originate from the same video frame, while pushing them apart otherwise. As shown in Table 2 (bottom), while the contrastive loss enhances LoCoNet’s in-domain performance by 0.3 mAP, it offers limited gains in out-of-domain scenarios. Interestingly, when combined with LASER, our method achieves an additional 1.4 and 1.0 mAP improvement on the Talkies and ASW benchmarks, respectively. These results highlight both the effectiveness and general applicability of LASER.
Models LASER-bench
Low Noise High Noise
TalkNet [tao2021someone] 94.8 77.8
TalkNet + LASER 95.0 82.1
LoCoNet [wang2024loconet] 96.2 86.7
LoCoNet + LASER 96.4 90.0
Table 3: Results on LASER-bench Combining LASER with TalkNet [tao2021someone] and LocoNet [tao2021someone] consistently improves performance on LASER-bench, with larger gains observed in scenarios with higher background noise.
5.3 Evaluation on LASER-bench
We further evaluate LASER on LASER-bench, treating it as an out-of-domain testbed for assessing robustness under real-world acoustic challenges. Models are trained on the AVA training set—composed primarily of clean, scripted movie scenes—and evaluated on the low-noise and high-noise subsets of LASER-bench, which feature modern online content with diverse and realistic background noise. This setup isolates the impact of acoustic variation without introducing major shifts in speaking style or visual context.
As shown in Table 3, LASER augmentation yields consistent gains: TalkNet’s mAP improves from 94.8 to 95.0 in low-noise conditions and from 77.8 to 82.1 in high-noise, a substantial improvement of 4.3 points. Similarly, LoCoNet improves from 96.2 to 96.4 in low-noise and from 86.7 to 90.0 in high-noise, gaining 3.3 points. Although both models are trained solely on AVA, LASER models generalize well to modern, noisy videos, with the largest gains observed under the most challenging acoustic conditions.
LASER-bench
@@ 5.3 Evaluation on LASER-bench
Trained with Lip Track Encoding (LTE) Consistency Loss Eval. with LTE AVA-Val Low Noise High Noise
N/A 95.2 96.2 86.7
95.1 96.1 85.7
95.3 96.5 90.0
95.3 96.4 90.0
Table 4: Importance of Each Component. Training solely with Lip Track Encoding (LTE) leads to a slight decrease in mAP on both AVA-val [roth2020ava] and LASER-bench, primarily due to the unavailability of lips landmarks on some face tracks. The consistency loss addresses this limitation, enhancing the model’s robustness.
5.4 Evaluation with Unsynchronized Audios
Model AVA Talkies ASW
LoCoNet [wang2024loconet] 75.9 72.1 56.7
LoCoNet + LASER 77.6 73.9 59.9
Table 5: Evaluation with swapped audio track. LASER significantly outperforms LoCoNet baseline when audio and visual frames are misaligned.
To test robustness against audio-visual misalignment, we simulate non-speaking scenarios by altering the audio stream while keeping the visual input unchanged. This reflects real-world situations like web conferencing, where off-screen or muted speakers may cause confusion. We consider two protocols on the AVA dataset: (1) audio-swapping, where the test audio is replaced with that from another video, and (2) audio-shifting, where a fixed delay is introduced. All experiments are conducted on the AVA test set, with accuracy computed as the average across all frames.
As shown in Figure 3 and Table 5, LASER consistently outperforms LoCoNet across both protocols. On audio-swapped AVA [roth2020ava], Talkies [alcazar2021maas], and ASW [kim2021look], it improves accuracy by 1.7%, 1.8%, and 3.2%, respectively. On audio-shifted AVA, it achieves at least a 2% gain, and maintains strong performance across all temporal shifts on Talkies and ASW. These results highlight LASER’s robustness under unsynchronized audio conditions.
@@ 5.5 Ablation Studies
We next present ablation studies of LASER. All ablation studies are conducted with the LoCoNet framework [wang2024loconet] unless otherwise noted.
5.5.1 Alternative Facial Landmark Integration
Model AVA Talkies ASW
LoCoNet [wang2024loconet] 95.2 88.4 88.4
w/ LP [Wang2019landmark_pooling] § 94.7 88.3 88.5
w/ LDI [Geeroms2022_2d_encoding] § 94.9 87.9 87.9
w/ LASER 95.3 89.0 89.5
Table 6: Comparison of different ways to incorporate lip landmarks. § Results are re-implemented in LoCoNet framework. LASER achieves the best performance over other alternative designs on both in-domain and out-of-domain evaluations.
Recall that we design a new encoding function to convert lip landmarks into 2D spatial features so that they can be integrated to visual features (Section 3.2.1). In this section, we compare our approach with other alternative designs to incorporate lip landmark information in to models.
Landmark Pooling (LP). Wang et al [Wang2019landmark_pooling] proposes to pool the visual features with lip landmark coordinates and use the resulting features in combination with original visual features for further processing.
Landmark as Discrete Inputs (LDI). Geeroms et al. [Geeroms2022_2d_encoding] directly uses convolution layers to process the raw landmark coordinates. A lip landmark track $f\in\mathbb{R}^{T\times N\times 2}$, where $T$ denotes the number of frames, $N$ is the number of landmarks, and 2 is the coordinate $(x,y)$, is convolved with convolution layers along the temporal dimension.
We implement these alternative designs in LoCoNet [wang2024loconet] and compare them with our approach. As shown in Table 6, LASER achieves the best performance over these baseline designs on both AVA and LASER-bench, highlighting its efficacy. Moreover, LASER only introduces negligible extra parameters to process lip landmarks compared to other approaches (details in Appendix).
5.5.2 Importance of Each Component
The lip landmark encoding and consistency loss are central to LASER, enabling the integration of lip information during training while addressing missing lip landmarks at test time. In practice, our landmark detector fails on 15% of AVA and 39% of LASER-bench test face tracks. As shown in Table 4, using lip landmark encoding alone leads to a slight mAP drop due to missing test-time landmarks. In contrast, adding the consistency loss allows the model to retain strong performance without relying on lip landmarks, showing that the visual encoder learns to capture lip motion implicitly. This removes the need for lip landmarks during inference and reduces latency.
Figure 4: Qualitative Comparison on LASER-bench. a) Ground-truth annotation. b) Detection results from our method LASER. c) Detection Results from baseline LoCoNet. Red: not speaking; Green: speaking; orange: incorrect predictions.
5.6 Qualitative Results
As shown in Figure 4, we present a qualitative comparison on LASER-Bench to illustrate how LASER handles high background noise (e.g., crowd cheering) more effectively than LoCoNet. In this clip, intense stadium noise overwhelms the audio signal, causing LoCoNet to miss genuine speaking events and produce false positives. In contrast, our method consistently localizes the true active speaker. This results in accurate speaking and non-speaking decisions where LoCoNet fails, highlighting the robustness of LASER in acoustically challenging, real-world scenarios.
6 Conclusion
In this work, we introduced LASER for active speaker detection. Unlike prior methods that rely on raw facial appearance, LASER explicitly guides the model to focus on lip landmarks, improving its ability to detect active speakers in challenging conditions. To bridge the gap in evaluating ASD performance under high background noise, we also introduced LASER-bench (LASER-Bench), a curated dataset featuring real-world videos with varying noise levels. Both quantitative and qualitative results show that LASER outperforms state-of-the-art models and achieves more reliable performance on LASER-Bench, demonstrating its robustness in noisy, unconstrained environments.
Acknowledgment
This work was supported in part by NSF IIS2404180, and Institute of Information & communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration) and (No. RS-2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training).
A Appendix
A.1 Training Details
We implement our approach LASER in three different state-of-the-art ASD frameworks, LoCoNet [wang2024loconet], TalkNet [tao2021someone], and Light-ASD [liao2023light]. We closely follow the training details for fair comparison. Specifically, for LoCoNet-based training, we use batch size of 4 and sample 200 frames per training example. Each model is trained with 25 epochs on 4 RTX 2080 GPUs. For Light-ASD and TalkNet, we use a batch size that contains at most 2000 frames and each model is trained on one A40 GPU. We use Adam [diederik2014adam] as our optimizer with learning rate of $5\times 10^{-5}$ that reduces by 0.5% per epoch. For LoCoNet and TalkNet, we set $\lambda_{av}=1$, $\lambda_{a}=0.4$, and $\lambda_{v}=0.4$; and for Light-ASD, we set $\lambda_{av}=1$ and $\lambda_{v}=0.5$, following the original work. Random resizing, cropping, horizontal flipping, and rotations are used as visual data augmentation operations and a randomly selected audio signal from the training set is added as background noise to the target audio [tao2021someone].
A.2 Implementation of Landmark Pooling and 2D Lip Encoding methods
In this section, we introduce implementation details of the Landmark Pooling method [Wang2019landmark_pooling] and Landmark as Discrete Inputs [Geeroms2022_2d_encoding]. Since these methods were originally proposed to work with simple CNN architectures, we adapt them to the state-of-the-art LoCoNet framework for fair comparison.
Figure A: Our implementation of Landmark Pooling [Wang2019landmark_pooling].
A.2.1 Landmark Pooling
Given a face track and the lips landmark detection results, we first obtain feature map $\mathcal{F}_{v}\in\mathbb{R}^{T\times C\times H\times W}$ after the first layer the ResNet, where $T$ denotes the number of frames, $H$ and $W$ are the spatial dimension of $f_{v}$, and $C$ is the number of channels, we follow [Wang2019landmark_pooling] to pool the feature with normalized lip landmark coordinates $\{(x_{i},y_{i})\}_{i=1}^{N}$ by $\mathcal{F}_{v}[x_{i},y_{i}]\in\mathbb{R}^{T\times C}$. Then, we concatenate all $N$ landmarks into $\mathcal{F}_{ld}\in\mathbb{R}^{T\times NC}$ and feed $\mathcal{F}_{ld}$ through a fully connected layer to get $\mathcal{F}^{\prime}_{ld}\in\mathbb{R}^{T\times D}$ where $D$ is the hidden dimension after the average pooling layer of ResNet. Finally, $\mathcal{F}^{\prime}_{ld}$ is concatenated to $\mathcal{F}^{\prime}_{v}$, the semantic feature output by ResNet and reduce the dimension with another fully-connected layer to obtain the final frame-level representation $\mathcal{F}\in\mathbb{R}^{T\times D}$. $\mathcal{F}$ is further processed by V-TCN and Context Modeling models along with audio feature to make prediction. The visualization of the method is presented in Figure A.
Figure B: Our implementation of 2D Lip Encoding method [Geeroms2022_2d_encoding].
A.2.2 Landmark as Discrete Inputs
Given a face track and the lips landmark detection results $f\in\mathbb{R}^{T\times N\times 2}$ where $T$ is the number of frames, $N$ is the number of landmarks, and 2 is the coordinate $(x,y)$. We adopt the architecture in [Geeroms2022_2d_encoding] to process $f$ with 3x3 convolution layers and convert the raw landmark coordinates into embeddings. The resulting embedding is flattened and concatednated to the visual feature in the same way of Landmark Pooling. The visualization of the architecture is presented in Figure B.
A.3 Parameter Effiency of LASER
As illustrated in Figure C, our proposed LASER approach not only achieves better performance but also only requires the negligible additional parameters compared to Landmark Pooling and Landmark as Discrete Inputs. This indicates that LASER’s design is significantly more parameter-efficient, which translates to reduced memory usage and faster training/inference. By minimizing the number of additional parameters without compromising accuracy, LASER offers a compelling balance between computational cost and performance gains.
Figure C: Number of additional parameters of each method. The y-axis is in log scale, and the number on each bar represents the true additional number of parameters. The x-axis contains 3 methods: LASER, Landmark Pooling [Wang2019landmark_pooling], and Landmark as Discrete Inputs [Geeroms2022_2d_encoding].
A.4 Hyperparameters Selection
We present our ablation study on hyperparamete selection, including the ablation on consistency loss weight and the Lip Track Encoding injection layer. The hyperparameter selection process is conducted through a holdout evaluation protocol (different from all protocols in main evaluation) by reversing the audio in original video clip to establish a non-speaking scenario. This avoids the risk of tuning on the test set and ensures generalizability.
A.4.1 ResNet vs 3D CNN